Will you please help me solve this problem? i have tried, but i just cannot figure out how to do it. It is on my semester exam study guide. here is the problem: John is 3 times as old as Beth. The sum of their ages is 36. How old is John.
i have put it in a table to clean my work up, and i made john equal X. and then i got to X=3x. i cannot figure out the rest. your help would be appreciated. thank you.

Answer by scott8148(6628)

"John is 3 times as old as Beth" ___ j = 3b

"The sum of their ages is 36" ___ j + b = 36 ___ 3j + 3b = 108

substituting j for 3b ___ 3j + (j) = 108 ___ 4j = 108